# If x^2+y^2=24 and xy=7 than solve for (x+y)^2
Solution ::
We know that
(x+y)^2 = x^2 + 2xy + y^2
thErEforE , wE gEt
(x+y)^2 = 24 + 2*7 = 38

Answer by Abdul Hannan(9)

# If x^2+y^2=24 and xy=7 than solve for (x+y)^2
Solution ::
We know that
(x+y)^2 = x^2 + 2xy + y^2
thErEforE , wE gEt
(x+y)^2 = 24 + 2*7 = 38